if you get paid 7.50 an hour for 40 hours a week and get 1.50 times your regular pay if you work over time how much would you get in 42 hours?
Answer by jim_thompson5910(35256)

Note: I'm assuming overtime is anything over 40 hours.


For the 40 hours you work, you earn a regular wage of $7.50, so you earn 7.50*40 = 300 dollars so far.


Then for the 2 additional overtime hours, you earn 1.50*7.50 = 11.25 dollars an hour. So you earn 2*11.25 = 22.50 dollars for those 2 extra hours.


In total, you have earned 300+22.50 = $322.50 for the 42 hours worked at $7.50 an hour.